# Weakly Supervised Object Segmentation by Background Conditional Divergence

### v1

@@ 3.1 Problem Formulation
[…]
@@ 3.2 Statistical Divergence Choice
The choice of the divergence $\mathit{D}$ will necessarily impact performance, computational cost, and the learning dynamics. To connect our approach to supervised segmentation tasks, the Wasserstein-2 distance is natural as it seeks a joint distribution $\mathbb{P}_{\tilde{X}_{c}X_{c}}$ that pairs the counterfactual $\tilde{X}_{c}$ and composite image $X_{c}$ to be as close as possible in a root-mean squared error sense
$\mathrm{W}_{2}(\mathbb{P}_{\tilde{X}_{c,\theta}},\mathbb{P}_{X|C_{b}=c})=\inf_{\mathbb{P}_{\tilde{X}_{c}X_{c}}\in\mathcal{P}}\sqrt{\mathbb{E}_{(\tilde{X}_{c},X_{c})\sim\mathbb{P}_{\tilde{X}_{c}X_{c}}}\lVert\tilde{X}_{c}-X_{c}\rVert_{2}^{2}},$ (9)
where $\mathcal{P}$ is the set of joint distributions with marginals equal to $\mathbb{P}_{\tilde{X}_{c,\theta}}$ and $\mathbb{P}_{X|C_{b}=c}$. The ideal joint distribution would pair the object in $X_{c}$ with a corresponding object in $X_{\neg c}$ and the background of $X_{c}$ with a corresponding background $R_{c}$ such that $\tilde{X}_{c}$ would match $X_{c}$ if the masking is correct. However, if the masking network is wrong the optimal joint distribution may create non-ideal pairings to compensate, possibly leading to local minima. Furthermore, estimating the Wasserstein distance in this primal form is challenging. Based on a review of divergences, including alternatives for the Wasserstein distance, in Appendix B, we select $\mathit{D}$ to be the energy-based sliced Wasserstein (EBSW) distance (Nguyen & Ho, 2023), which avoids adversarial training of discriminators required for the variational estimates of $f$-divergences in GANs (Goodfellow et al., 2014; Nowozin et al., 2016) and has desirable computational and statistical characteristics—in particular it does not suffer from curse of dimensionality (Nguyen & Ho, 2023)—making it broadly applicable without many hyper-parameters. For the hyper-parameters of the estimate of the EBSW, described by equation 14 in Appendix B, we choose $p=2$ to match equation 9, $f$ to be the exponential function, and $L=1000$ slices.
@@ 3.3 Background Clustering
For clustering the backgrounds, we propose using a two-stage process of either an autoencoder (AE) or an encoder trained with self-supervised contrastive learning (CL) followed by K-means on the encoders output. The positive pair in the CL setting is defined as a randomly cropped background image (no less than 80%) and resized to match the original image.
@@
The encoder training and subsequent cluster assignment can be applied to all or a subset of the background-only images. Then, during the training of the masking network background-only images and composite images (or mostly background images from nearby crops) are assigned based on the nearest cluster to their encoding, yielding the label $C_{b}$.
3.4 Data
@@
While our goal is to perform real-world weakly supervised object segmentation on sonar images, datasets with ground truth labels are limited. We developed and initially validated our proposed method on artificial single-channel data constructed from combining real-world backgrounds—either textures (photographic images of materials with various textures (Brodatz, 1966) under controlled lighting) or seafloor images obtained from sonar (Cobb & Zare, 2014)—with objects from standard object classification datasets: standard MNIST, FashionMNIST (Xiao et al., 2017), and dSprites (Matthey et al., 2017). On these toy datasets we also ensure that the object is independently chosen from the background. The description and results on these artificial datasets is included in Appendix C.1.
@@
We validate our approach on purely real datasets. The first dataset is a real side-scan sonar (SSS) dataset (Sethuraman et al., 2024), which contains a variety of foreground objects corresponding to different sites that contain shipwrecks in Lake Michigan along with ground truth segmentation masks for the shipwrecks in the seabed. The second dataset, is the Caltech-UCSD Birds-200-2011 (CUB-200-211) dataset (Wah et al., 2011), which consists of three-channel (RGB) color images of 200 species of birds on various backgrounds with segmentations.
3.4.1 Side Scan Sonar with Shipwreck Objects
@@
The first dataset is a real side-scan sonar (SSS) dataset (Sethuraman et al., 2024), which contains a variety of foreground objects corresponding to different sites that contain shipwrecks in Lake Michigan along with ground truth segmentation masks for the shipwrecks in the seabed. The SSS dataset (Sethuraman et al., 2024) contains 28 distinct shipwrecks totaling 286 high-resolution images with segmentations. The dataset is already divided 50/50 into training and test respecting the variety of sites. We further split the training data into $70\%$ training and $30\%$ validation.
@@
For each sonar scene, we crop out the black strip in the middle (i.e., the blind spot directly below the sonar platform) and we flip the right side of the sonar image so that range increases consistently from right to left in images and the acoustic shadows all have the same orientation. Then, we crop 500 $128\times 128$ images from the 143 images in the training set. We normalize each crop by its mean and then clip any value larger than 16. Then, we resample each crop to $64\times 64$. We use ground truth labels to label each crop as composite or background-only. Although this approach uses supervised information, the idea of weak supervision is that each crop could be quickly manually labeled or one could use a statistical anomaly detection algorithm for the presence of objects in groups. In either case, these are beyond the scope of this work.
@@
3.4.2 CUB-200-211
@@
To illustrate the broader applicability of our approach across diverse imaging modalities, we also evaluate it on a fourth dataset comprising natural images from the CUB dataset. We also demonstrate our work on the CUB-200-211 dataset (Wah et al., 2011), which consists of 11,788 images of birds from 200 species on various backgrounds. We chose this dataset to demonstrate that our work can also be applied to natural images, although the assumptions of independence of object and background are likely violated for birds with different habitats. Each image is min-max normalized such that all pixel values are between $[0,1]$. We report the results for the separate testing set which consists of the half of the images, specifically, the train-test split has 5994 train and 5794 test images. We further split the training data into $70\%$ training and $30\%$ validation. Additionally, when benchmarking against other unsupervised object segmentation approaches, we follow Chen et al. (2019) and use their split consists of 10,000 images for training, 1,000 for testing, and the remaining 778 for validation.
3.5 Implementation Details
@@
The AE or CL encoder (CLE) and K-means are trained on background-only images. The AE architecture is composed of 4 convolution layers each followed by a batch normalization layer and ReLU activation. The number of kernels in each layer is $\{64,128,256,512\}$. A fully connected bottleneck layer of dimension $o=20$ is used to represent the latent space. Similarly, the decoder is composed of one fully connected layer and 4 transpose convolutions with the same number of kernels as the encoder but with a reversed order. We add a sigmoid layer at the end of the decoder as we find it accelerates training the AE. We used the Adam optimizer to optimize the AE, with a fixed learning rate of $0.0001$, and batch size of $4096$. The CLE follow the exact design of the encoder part in the AE but we increased the latent dimension to $o=200$ for CUB dataset.
@@
For the SSS dataset, a nearby patch of seafloor without any object can be passed to the AE/CLE + Sinkhorn K-means to infer the background of a foreground object image. The entropic regularization parameter in the Sinkhorn algorithm was set to 0.01. For CUB-200-2011, we train CLE and K-means and we use patches of size $25\%$ of the image width and height taken from the composite images as background-only images: during CLE training, we use the ground truth masks to ensure that the patch does not contain any objects, but during training of the masking network we use a random patch cropped from the input image. In this case, we use the ground truth to satisfy the assumption that representative background-only images are available; however we train the masking network without any ground truth knowledge. For the contrastive learning, we use augmentation techniques such as random cropping to assign positive pairs. We also used random horizontal flipping for the CUB dataset.
@@
For SSS dataset, we choose clusters from the set $\{2,3,4,5\}$ but we found that they perform very similarly so we only report the 4 clusters results. For CUB-200-2011, we train models with varying the number of clusters, choosing values from the set $\{5,10,15\}$. In both cases, the choice of cluster numbers is guided by the visual distinctiveness of example images in each cluster: we increase the number of clusters until further subdivision yields negligible improvement in visual separation.
@@
We choose our segmentation model to be the U-Net (Ronneberger et al., 2015b) with default parameters. We choose D-adaptation optimizer to optimize U-Net (Defazio & Mishchenko, 2023). The learning rate for the U-Net is $1$ and the batch size is $400$. We selected the largest batch size the GPU we used for training could handle. For the SSS and CUB datasets, we run the training for $150$ and $200$ epochs, respectively. We choose the model with lowest validation loss $\mathcal{L}$ shown in equation 8. We implement our approach using the PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) framework. Only one GPU, a Tesla V100-SXM2-32GB, is used for training.

### v2

@@ 3.1 Problem Formulation
$\min_{\theta\in\Theta}\mathcal{L}_{\mathit{D}}(\theta)+\mathcal{L}_{\text{Bg}}(\theta),$ (8)
@@
In practice, the training data are $\{(x_{i},c_{i})\}_{i=1}^{N_{\text{Comp.}}}\subset\mathcal{X}\times\{1,\ldots,K\}$ and $\{(r_{i},c^{\prime}_{i})\}_{i=1}^{N_{\text{Bg}}}\subset\mathcal{R}\times\{1,\ldots,K\}$, where $N_{\text{Comp.}}$ and $N_{\text{Bg}}$ are the total number of composite and background-only images, respectively. We define the empirical conditional distributions $\mathbb{P}_{X|C_{b}=c}=\frac{1}{|\{i|c_{i}=c\}|}\sum_{i|c_{i}=c}\delta_{x_{i}}$, $\mathbb{P}_{X|C_{b}\neq c}=\frac{1}{|\{i|c_{i}\neq c\}|}\sum_{i|c_{i}\neq c}\delta_{x_{i}}$, and $\mathbb{P}_{R|C_{b}=c}=\frac{1}{|\{i|c^{\prime}_{i}=c\}|}\sum_{i|c^{\prime}_{i}=c}\delta_{r_{i}}$, where $\delta_{x}$ is the Dirac measure $\delta_{x}(\mathcal{A})=\begin{cases}1&x\in\mathcal{A}\\ 0&x\notin\mathcal{A}\end{cases},\quad\forall\mathcal{A}\subseteq\mathcal{I}$. For a given background $c$, three sets of mini-batches of size $B$ are sampled $\{x_{\neg c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{X|C_{b}\neq c}$, $\{r_{c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{R|C_{b}=c}$, $\{x_{c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{X|C_{b}=c}$, which correspond to $X_{c}$, $X_{\neg c}$, and $R_{c}$, respectively. Then the counterfactuals are produced $\tilde{x}_{c}^{(i)}=(1-M_{\theta}(x_{\neg c}^{(i)}))\circ r_{c}^{(i)}+M_{\theta}(x_{\neg c}^{(i)})\circ x_{\neg c}^{(i)},\quad i\in\{1,\ldots,B\}$. Finally, the divergence $D(\hat{\mathbb{P}}_{\tilde{X}_{c,\theta}},\;\hat{\mathbb{P}}_{X|C_{b}=c})$ is computed from the empirical distributions $\hat{\mathbb{P}}_{\tilde{X}_{c,\theta}}=\tfrac{1}{B}\sum_{i=1}^{B}\delta_{\tilde{x}_{c}^{(i)}}$ and $\hat{\mathbb{P}}_{X|C_{b}=c}=\tfrac{1}{B}\sum_{i=1}^{B}\delta_{x_{c}^{(i)}}$. Denoting the mini-batch loss functions as $\hat{\mathcal{L}}$, we summarize the training procedure in Algorithm 1.
@@
Algorithm 1 Training Procedure
1:Model $M_{\theta}$, distributions $\mathbb{P}_{X|C_{b}=c}$, $\mathbb{P}_{X|C_{b}\neq c}$, and $\mathbb{P}_{R|C_{b}=c}$ for $c\in\{1,\ldots,K\}$, batch size $B$
2:for each training iteration do
3:  for each category $c\in\{1,\ldots,K\}$ do
4:   Sample mini-batches: $\{x_{\neg c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{X|C_{b}\neq c},\quad\{r_{c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{R|C_{b}=c},\quad\{x_{c}^{(i)}\}_{i=1}^{B}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathbb{P}_{X|C_{b}=c}.$
5:   $\tilde{x}_{c}^{(i)}=(1-M_{\theta}(x_{\neg c}^{(i)}))\circ r_{c}^{(i)}+M_{\theta}(x_{\neg c}^{(i)})\circ x_{\neg c}^{(i)},\quad i\in\{1,\ldots,B\}$
6:   $\hat{\mathcal{L}}^{(c)}_{D}(\theta)=D(\hat{\mathbb{P}}_{\tilde{X}_{c,\theta}},\;\hat{\mathbb{P}}_{X|C_{b}=c}),\quad\hat{\mathbb{P}}_{\tilde{X}_{c,\theta}}=\tfrac{1}{B}\sum_{i=1}^{B}\delta_{\tilde{x}_{c}^{(i)}},\quad\hat{\mathbb{P}}_{X|C_{b}=c}=\tfrac{1}{B}\sum_{i=1}^{B}\delta_{x_{c}^{(i)}}$.
7:   $\hat{\mathcal{L}}^{(c)}_{\text{Bg}}(\theta)=\tfrac{1}{B}\sum_{i=1}^{B}\|M_{\theta}(r_{c}^{(i)})\|_{F}^{2}$ (Background loss)
8:  end for
9:  Update $\theta$ using gradient step on the total loss: $\hat{\mathcal{L}}(\theta)=\sum_{c=1}^{K}\left(\hat{\mathcal{L}}^{(c)}_{D}(\theta)+\hat{\mathcal{L}}^{(c)}_{\text{Bg}}(\theta)\right)=\hat{\mathcal{L}}_{D}(\theta)+\hat{\mathcal{L}}_{\text{Bg}}(\theta)$.
10:end for
@@ 3.2 Statistical Divergence Choice
The choice of the divergence $\mathit{D}$ will necessarily impact performance, computational cost, and the learning dynamics. To connect our approach to supervised segmentation tasks, the Wasserstein-2 distance is natural as it seeks a joint distribution $\mathbb{P}_{\tilde{X}_{c,\theta}X_{c}}$ that pairs the counterfactual $\tilde{X}_{c}$ and composite image $X_{c}$ to be as close as possible in a root-mean squared error sense
$\mathrm{W}_{2}(\mathbb{P}_{\tilde{X}_{c,\theta}},\mathbb{P}_{X|C_{b}=c})=\inf_{\mathbb{P}_{\tilde{X}_{c,\theta}X_{c}}\in\mathcal{P}}\sqrt{\mathbb{E}_{(\tilde{X}_{c},X_{c})\sim\mathbb{P}_{\tilde{X}_{c,\theta}X_{c}}}\lVert\tilde{X}_{c}-X_{c}\rVert_{2}^{2}},$ (9)
where $\mathcal{P}$ is the set of joint distributions with marginals equal to $\mathbb{P}_{\tilde{X}_{c,\theta}}$ and $\mathbb{P}_{X|C_{b}=c}$. The ideal joint distribution would pair the object in $X_{c}$ with a corresponding object in $X_{\neg c}$ and the background of $X_{c}$ with a corresponding background $R_{c}$ such that $\tilde{X}_{c}$ would match $X_{c}$ if the masking is correct. However, if the masking network is wrong, the optimal joint distribution may create non-ideal pairings to compensate, possibly leading to local minima.
Estimating the Wasserstein distance in this primal form is challenging. Based on a review of divergences, including alternatives for the Wasserstein distance, in Appendix B, we select $\mathit{D}$ to be the energy-based sliced Wasserstein (EBSW) distance (Nguyen & Ho, 2023), which avoids adversarial training of discriminators required for the variational estimates of $f$-divergences in GANs (Goodfellow et al., 2014; Nowozin et al., 2016) and has desirable computational and statistical characteristics. In particular it does not suffer from the curse of dimensionality (Nguyen & Ho, 2023), making it broadly applicable without many hyperparameters. For the hyperparameters of the estimate of the EBSW, described by equation 14 in Appendix B, we choose $p=2$ to match equation 9, $f$ to be the exponential function, and $L=1000$ slices.
@@ 3.3 Background Clustering
For clustering the backgrounds, we propose using a two-stage process of either an autoencoder (AE) or an encoder trained with self-supervised contrastive learning (CL) followed by K-means on the encoder’s output. The positive pair in the CL setting is defined as a randomly cropped background image (no less than 80%) and resized to match the original image.
@@
The encoder training and subsequent cluster assignment can be applied to all or a subset of the background-only images. Then, during the training of the masking network background-only images and composite images (or mostly background images from nearby crops) are assigned based on the nearest cluster to their encoding, yielding the label $C_{b}$. This produces the background-labeled images underlying the distributions required for Algorithm 1.
4 Data
@@
While our goal is to perform real-world weakly supervised object segmentation on sonar images, datasets with ground truth labels are limited. We developed and initially validated our proposed method on artificial single-channel data constructed from combining real-world backgrounds—either textures (photographic images of materials with various textures (Brodatz, 1966) under controlled lighting) or seafloor images obtained from sonar (Cobb & Zare, 2014)—with objects from standard object classification datasets: standard MNIST, FashionMNIST (Xiao et al., 2017), and dSprites (Matthey et al., 2017). On these toy datasets we also ensure that the object is independently chosen from the background. Details and specific implementation of each of the toy datasets (Textures+MNIST; Textures+FashionMNIST; Textures+dSprites; and SAS+dSprites) are included in Appendix C.1.
@@
We validate our approach on purely real datasets. The first is the AI4Shipwrecks dataset consisting of side-scan sonar (SSS) images (Sethuraman et al., 2024), which contain a variety of foreground objects corresponding to different sites that contain shipwrecks in Lake Michigan along with ground truth segmentation masks for the shipwrecks in the seabed. The second is the SAS-Clutter dataset consisting of small image snippets taken from high-frequency synthetic aperture sonar (SAS) collected by Naval Surface Warfare Center Panama City Division (NSWC PCD). The seafloor clutter objects in the images vary in size from fractions of a meter to several meters in the longest dimension. The third is the Caltech-UCSD Birds-200-2011 (CUB-200-211) dataset (Wah et al., 2011), which consists of three-channel (RGB) color images of 200 species of birds on various backgrounds with segmentations.
4.1 AI4Shipwrecks
@@
The AI4Shipwrecks dataset (Sethuraman et al., 2024) consists of 286 high-resolution SSS images with a variety of foreground objects corresponding to 28 distinct sites that contain shipwrecks in Lake Michigan along with ground truth segmentation masks for the shipwrecks in the seabed. The dataset is already divided 50/50 into training and test respecting the variety of sites. We further split the training data into $70\%$ training and $30\%$ validation.
@@
For each sonar scene, we crop out the black strip in the middle (i.e., the blind spot directly below the sonar platform), and we flip the right side of the sonar image so that range increases consistently from right to left in images and the acoustic shadows all have the same orientation. Then, we crop 500 $128\times 128$ images from each of the 143 images in the training set, creating 71,500 crops. As sonar data can have an extremely large dynamic range, we normalize each crop by dividing by its mean pixel value and then clipping any value larger than 16. We then resample each crop to $64\times 64$. We use ground truth labels to label each crop as composite or background-only. Although this approach uses supervised information, the idea of weak supervision is that each crop could be quickly manually labeled, or one could use a statistical anomaly detection algorithm for the presence of objects in groups.
@@
4.2 SAS-Clutter
It is often necessary to identify and analyze small seafloor objects within SAS images. The SAS-Clutter dataset consists of approximately $10,000$ sonar images of underwater seafloor clutter objects collected using a high-frequency SAS from various coastal locations. The object image snippets were cropped from larger seafloor image survey data by first running a Reed-Xialoi (RX) based anomaly detector on the seafloor data and then manually selecting the snippets that actually contain objects (as opposed to other anomalies such as fish clouds, interference, or random noise). Seafloors are typically sparse, with very few objects, and infrequently change in composition/texture from one location to another. Thus, we make the assumption that crops taken from seafloor data adjacent to the selected objects will both be empty of objects and contain the same type of seafloor as its corresponding object. For each object image, we randomly selected a non-overlapping adjacent patch of seafloor to generate the background-only data. A validation set was created by manually drawing masks on a small random subset of approximately $200$ objects. Furthermore, a set of approximately $2000$ objects was held out for testing. Each object image was normalized by thresholding the pixel values at $16$ times its mean pixel value and then rescaling to $[0,1]$. Lastly, each image was resized to $64\times 64$ using bilinear interpolation.
4.3 CUB-200-211
@@
To illustrate the broader applicability of our approach across diverse imaging modalities, we also evaluate it on a dataset comprising natural images from the CUB-200-211 dataset (Wah et al., 2011), which consists of 11,788 images of birds from 200 species on various backgrounds. We chose this dataset to demonstrate that our work can also be applied to natural images, although the assumptions of independence of object and background are likely violated for birds with different habitats. Each image is min-max normalized such that all pixel values are between $[0,1]$. We report the results for the separate testing set which consists of roughly half of the images; specifically, there are 5994 train images and 5794 test images. We further split the training data into $70\%$ training and $30\%$ validation. Additionally, when benchmarking against other unsupervised object segmentation approaches, we follow Chen et al. (2019) and use their split, which consists of 10000 images for training, 1000 for testing, and the remaining 778 for validation.
5 Implementation Details
@@
The AE or CL encoder (CLE) and K-means are trained on background-only images. The AE architecture is composed of 4 convolution layers each followed by a batch normalization layer and ReLU activation. The number of kernels in each layer are $64,128,256,512$, respectively. A fully connected bottleneck layer of dimension $o=20$ is used to represent the latent space. Similarly, the decoder is composed of one fully connected layer and 4 transpose convolutions with the same number of kernels as the encoder but in reversed order. We add a sigmoid layer at the end of the decoder. We use the Adam optimizer to optimize the AE, with a fixed learning rate of $0.0001$, and batch size of $4096$. The CLE follow the exact design of the encoder part in the AE but we increased the latent dimension to $o=200$ for CUB dataset.
@@
For the AI4Shipwrecks and SAS-Clutter datasets, a nearby patch of seafloor without any object can be passed to the AE/CLE + Sinkhorn K-means to infer the background of a foreground object image. The entropic regularization parameter in the Sinkhorn algorithm was set to 0.01. For CUB-200-2011, we train CLE and K-means and we use patches of size $25\%$ of the image width and height taken from the composite images as background-only images: during CLE training, we use the ground truth masks to ensure that the patch does not contain any objects, but during training of the masking network, we use a random patch cropped from the input image. In this case, we use the ground truth to satisfy the assumption that representative background-only images are available; however we train the masking network without any ground truth knowledge. For the contrastive learning, we use augmentation techniques such as random cropping to assign positive pairs. We also use random horizontal flipping for the CUB dataset.
@@
For AI4Shipwrecks and SAS-Clutter datasets, we choose clusters from the set $\{2,3,4,5\}$ but we found that they perform very similarly so we only report the 4 clusters results. For CUB-200-2011, we train models with varying the number of clusters, choosing values from the set $\{5,10,15\}$. In both cases, the choice of cluster numbers is guided by the visual distinctiveness of example images in each cluster: we increase the number of clusters until further subdivision yields negligible improvement in visual separation.
@@
We choose our segmentation model to be the U-Net architecture (Ronneberger et al., 2015b) with default hyperparameters. We choose D-adaptation optimizer to optimize U-Net (Defazio & Mishchenko, 2023). The learning rate for the U-Net is $1$ and the batch size is $400$. We selected the largest batch size the GPU used for training could handle. For the AI4Shipwrecks and CUB datasets, we run the training for $150$ and $200$ epochs, respectively. We choose the model with lowest validation loss $\mathcal{L}$ shown in equation 8. We implement our approach using the PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019) framework. Only one GPU, a Tesla V100-SXM2-32GB, is used for training.